OpenStudy (anonymous):

The height of an object that is dropped or thrown straight up or down is given by the quadratic function h(x) = –16x^2+ v0x + h0, where x is the time in seconds, v0 is the velocity at 0 seconds, and h0 is the height at 0 seconds. Suppose that a ball is thrown straight up from a height of 2 feet at a velocity of 31 feet per second.

OpenStudy (fifciol):

What is the question?

OpenStudy (anonymous):

what is the maximum height of the ball ? what time does the ball reach the maximum height ?

OpenStudy (fifciol):

ok, h0 is the initial height and v0 is the initial velocity. what does our equation look now when you put that numbers?

OpenStudy (anonymous):

h(x)=-16^2+31x+2

OpenStudy (anonymous):

i forgot the x in between 16 and 2

OpenStudy (fifciol):

yes, this is the equation of a parabola that looks like this:  When that function reaches the maximum? when h'(x)=0. Can you solve it?

OpenStudy (anonymous):

Yes ?

OpenStudy (fifciol):

What is derivative of that function?

OpenStudy (anonymous):

Is it -2 and 3

OpenStudy (fifciol):

are you familiar with calculus?

OpenStudy (anonymous):

No

OpenStudy (fifciol):

ok, do you know how to find the vertex of parabola? that would be your max height

OpenStudy (anonymous):

the vertex is 2 , right ?

OpenStudy (fifciol):

no it's \[-\frac{ \Delta}{ 4a }\]

OpenStudy (anonymous):

the quadratic formula ?

OpenStudy (fifciol):

or substitute x=-b/2a

OpenStudy (fifciol):

you should find hmax=31/32

OpenStudy (anonymous):

Yes i got x =31/32 or 1089/64

OpenStudy (fifciol):

you shouldget only 31/32

OpenStudy (anonymous):

yes

OpenStudy (fifciol):

then put that in place of h and solve for x. If you find two solutions choose positive one(time cannot be negative)

OpenStudy (anonymous):

so im replacing 2

OpenStudy (anonymous):

1/124